approximate the following logarithms using the properties of logarithms given logb2=0.3562,logb3=0.5646,logb5=0.8271:
logb2/5
logb5b
logb30

Answer by stanbon(48507)

approximate the following logarithms using the properties of logarithms given logb2=0.3562,logb3=0.5646,logb5=0.8271:
--------------------------------------------
logb(2/5) = logb(2)-logb(5)
--------------------------------
logb(5b) = logb(5) + logb(b) = logb(5) + 1
--------------------------------
logb(30) = logb[2*3*5] = logb(2) + logb(3) + logb(5)
---------------------------------
I'll let you do the substitutions and the arithmetic.
==================================
Cheers,
Stan H.